# Enhancing Interpretability in Deep Reinforcement Learning through Semantic Clustering

### v4

@@ 4.2 Semantic Clustering in DRL
Human Evaluation In addition to qualitatively analyzing the mean images, we hired 15 human evaluators to validate the semantic clustering properties. Specifically, video sequences from each episode are segmented into multiple clips based on the cluster each frame belongs to, and these clips are grouped by cluster for evaluators to review. Each evaluator watched these grouped clips and responded to three interpretability-related statements for two out of a set of three games (Jumper, Fruitbot, and Ninja). The response for each question was chosen from a five-point Likert scale with the following options: Strongly Disagree (1), Disagree (2), Neutral (3), Agree (4), and Strongly Agree (5). Further details on the human evaluation procedure are provided in Appendix I.
The statements and the results of the human evaluation are provided in subsection 4.2. The mean scores for all statement-environment combinations are greater than 4, with the exception of statements 1 and 3 for FruitBot, for which the lower bounds on the mean set by the standard error of the mean (SEM) are 3.99 and 3.92 respectively. The slightly lower score on FruitBot may be caused by the behavior description of clusters, the agent’s relative distance to the wall ahead (far/near) and the agent’s relative position on the screen (left/center/right) require a higher degree of subjective judgment. In contrast, Jumper has a clear radar for direction and position information, and Ninja has more explicit behavioral reference objects, e.g., ledges and mushrooms. Overall, these results suggest that humans generally agree that our model possesses semantic clustering properties and supports interpretability.
@@ 4.3 Model and Policy Analysis
Figure 5: Hover examples in the FDR space of Ninja. We observe a sub-cluster in the FDR space as an example from the overall perspective (a) and the zoomed-in perspective (b), (c), and (d). The agent is standing on the edge of a ledge. Although the scenarios of (b), (c), and (d) are different, the proposed method effectively clusters semantically consistent features together in the FDR space.
@@
Figure 6: Policy analysis examples in Ninja.
@@
We present policy analysis examples in Figure 6, leveraging clustering results from our method. Figures 6(a) and 6(b) illustrate consecutive states assigned to clusters 5 and 7, respectively. These examples show that the decision to approach the mushroom depends not on the mushroom’s presence but on detecting a right-side wall. This finding is confirmed in 6(c), another state from cluster 7 without any mushroom present, where the agent repeatedly moves along the ledge, incorrectly perceiving conditions for behaviors in cluster 7. Figures 6(d) and 6(e) depict states within the same episode: in 6(d), the agent initially plans to jump onto a lower ledge (cluster 5), but upon seeing a higher, safer ledge in 6(e), it shifts its strategy accordingly (cluster 3). Similarly, 6(f) shows a state that has been assigned to cluster 3 by our model, helping us anticipate the agent’s future behavior of jumping onto the higher ledge. These analyses demonstrate how our model helps clarify policy behaviors, uncover decision-making structures, and identify potential issues.
@@ 5 Limitations and Future Work
Our approach has several limitations to be addressed in future work. First, it relies on clear semantic distributions, which can become unstable when policies deviate significantly from optimal behavior, resulting in ambiguous clusters. More robust clustering methods may be needed to improve stability. Second, as the method is unsupervised, selecting an appropriate number of clusters is crucial—too few reduce clarity, while too many cause semantic fragmentation. We used eight clusters to balance interpretability and granularity, but future work could explore adaptive techniques that adjust the number of clusters based on task complexity. For further analysis of the impact of cluster numbers on performance and interpretability, see Appendix D. Furthermore, policy interpretations are manually described. In future work, we aim to automate behavior summarization and explanation. Lastly, we plan to extend this method to other DRL algorithms, benchmarks, and settings.
@@ Appendix A Architecture, Hyperparameters, and Computational Costs
In Equation 2 of the main paper, $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are 500 and 1, respectively. $f_{\text{control}}$ is updated every 50 iterations according to the following expression:
$f_{\text{control}}=\min\left(\frac{s_{\text{mean}}}{0.8\cdot s_{\text{highest}}},1\right),$ (A.1)
@@
All hyperparameters introduced in our method, except for the number of embeddings, were chosen through performance tuning to optimize the model’s overall performance. The number of embeddings in the vector quantizer was determined by ensuring that each cluster maintained a singular semantic interpretation.
@@ Appendix E Clustering Quality under Different Dimensionality‑Reduction Methods
To quantify how well various dimensionality‑reduction (DR) techniques preserve the semantic clusters discovered by our model, we project the same set of high‑dimensional state features—collected following the procedure described in subsection 4.1—using five popular DR baselines: UMAP, Isomap, PCA, MDS, and t‑SNE. These metrics were selected to evaluate both the compactness within clusters and the separation between clusters in the reduced feature space.
@@ Appendix F Semantic Formation in Clusters
To provide insight into the generation of semantic clustering in the feature space, we conducted the statistical analysis as follows. Initially, we collected 50,000 states using the trained model and computed the mean image for each cluster. We then calculated the mean and standard deviation of the pixel distances between the states in each cluster and their corresponding mean image. These measurements were averaged across the eight clusters. Additionally, we evaluated the probability of cluster transitions in these 50,000 states. We repeated the same statistical analysis on the 50,000 states using the untrained model, and the results are presented in Table 4. In the trained models, the pixel differences between the states within clusters increased, indicating the presence of states with similar semantics but larger pixel distances. Furthermore, the probability of cluster transitions decreased as it approached the required transition in the scene.
Table 4: Comparison metrics of the semantically trained model verse the raw model on various environments.
Model Cluster transition probability Pixel distance mean (Std. Dev.)
@@
Ninja Trained 0.2680 141.50 (67.88)

### v5

@@ 4.2 Semantic Clustering in DRL
Human Evaluation In addition to qualitatively analyzing the mean images, we hired 15 human evaluators to validate the semantic clustering properties. Evaluators were adults (18+), native or highly proficient English speakers, with basic video game experience and a brief training session. Specifically, video sequences from each episode are segmented into multiple clips based on the cluster each frame belongs to, and these clips are grouped by cluster for evaluators to review. Each evaluator watched these grouped clips and responded to three interpretability-related statements for two out of a set of three games (Jumper, Fruitbot, and Ninja). The response for each question was chosen from a five-point Likert scale with the options: Strongly Disagree (1), Disagree (2), Neutral (3), Agree (4), and Strongly Agree (5). Further details on the evaluation procedure are provided in Appendix I.
The statements and the results of the human evaluation are provided in § 4.2. The mean scores for all statement-environment combinations are greater than 4, with the exception of statements 1 and 3 for FruitBot, for which the lower bounds on the mean set by the standard error of the mean (SEM) are 3.99 and 3.92 respectively. The slightly lower score on FruitBot may be caused by the behavior description of clusters, the agent’s relative distance to the wall ahead (far/near) and the agent’s relative position on the screen (left/center/right) require a higher degree of subjective judgment. In contrast, Jumper has a clear radar for direction and position information, and Ninja has more explicit behavioral reference objects, e.g., ledges and mushrooms. Overall, these results suggest that humans generally agree that our model possesses semantic clustering properties and supports interpretability.
@@ 4.3 Model and Policy Analysis
Figure 5: Hover examples in the FDR space of Ninja. We observe a sub-cluster in the FDR space as an example from an zoomed-out perspective (a) and zoomed-in perspectives (b), (c), and (d). The agent is standing on the edge of a ledge. Although the scenarios of (b), (c), and (d) are different, the proposed method effectively clusters semantically consistent features together in the FDR space.
@@
Figure 6: Policy analysis examples in Ninja, showing states assigned to different clusters.
@@
We present policy analysis examples in Figure 6, leveraging clustering results from our method. Figures 6(a) and 6(b) show consecutive states assigned to clusters 5 and 7, respectively (see Figure 3 and Table 1). In 6(a), the right-side wall is absent and the agent walks along the ledge (cluster 5); in 6(b), a right-side wall appears and the agent transitions to cluster 7, approaching the mushroom. Because the mushroom is visible in both frames, this cluster change is driven by detecting the right-side wall rather than by the mushroom’s presence. This finding is confirmed in 6(c), which shows another state from cluster 7 without any mushroom present, where the agent continues along the ledge, incorrectly perceiving the conditions for cluster-7 behavior. Figures 6(d) and 6(e) depict states within the same episode: in 6(d), the agent initially plans to jump onto a lower ledge (cluster 5), but upon seeing a higher, safer ledge in 6(e), it shifts its strategy accordingly (cluster 3). Similarly, 6(f) shows a state that has been assigned to cluster 3 by our model, helping us anticipate the agent’s future behavior of jumping onto the higher ledge. These analyses demonstrate how our model helps clarify policy behaviors, uncover decision-making structures, and identify potential issues.
@@ 5 Limitations and Future Work
Our approach has several limitations to be addressed in future work. First, it relies on clear semantic distributions, which can become unstable when policies deviate significantly from optimal behavior, resulting in ambiguous clusters. More robust clustering methods may be needed to improve stability. Second, as the method is unsupervised, selecting an appropriate number of clusters is crucial—too few clusters reduces clarity, while too many clusters causes semantic fragmentation. We used eight clusters to balance interpretability and granularity, but future work could explore adaptive techniques that adjust the number of clusters based on task complexity, the elbow method, silhouette-score optimization, etc. For further analysis of the impact of cluster numbers on performance and interpretability, see Appendix D. Furthermore, policy interpretations are manually described. In future work, we aim to automate behavior summarization and explanation (e.g., using GPT-4V). In addition, since the FDR module optimizes pairwise similarities, a natural extension is to replace the current affinity with alternative measures (e.g., cosine similarity or bisimulation metrics). Lastly, we plan to extend this method to other DRL algorithms, benchmarks, and settings.
@@ 6 Conclusion
Acknowledgments
Research was sponsored by the Army Research Office and was accomplished under awards W911NF-20-1-0002 and W911NF-24-2-0034. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. We gratefully acknowledge the anonymous reviewers and the program committee for their thoughtful feedback and constructive suggestions, which helped improve this work. We also thank Huy Le and Robert Lopez for their valuable assistance with the human evaluation.
@@ Appendix A Architecture, Hyperparameters, and Computational Costs
In Equation 2 of the main paper, $w_{\text{FDR}}$ and $w_{\text{VQ-VAE}}$ are 500 and 1, respectively. $\lambda_{\text{ctrl}}$ is updated every 50 iterations according to the following expression:
$\lambda_{\text{ctrl}}=\min\left(\frac{s_{\text{mean}}}{0.8\cdot s_{\text{highest}}},1\right),$ (A.1)
@@
All hyperparameters introduced in our method, except for the number of embeddings, were chosen through performance tuning to optimize the model’s overall performance. The number of embeddings in the vector quantizer was determined by ensuring that each cluster maintained a singular semantic interpretation. During hyperparameter tuning, we found that performance is primarily influenced by $w_{\text{FDR}}$, $w_{\text{VQ-VAE}}$, and $\lambda_{\text{ctrl}}$, and is more robust to the number of VQ embeddings and the degrees of freedom in Equation 3.
@@ Appendix E Clustering Quality under Different Dimensionality‑Reduction Methods
To quantify how well various dimensionality‑reduction (DR) techniques preserve the semantic clusters discovered by our model, we project the same set of high‑dimensional state features—collected following the procedure described in § 4.1—using five popular DR baselines: UMAP, Isomap, PCA, MDS, and t‑SNE. These metrics were selected to evaluate both the compactness within clusters and the separation between clusters in the reduced feature space.
@@ Appendix F Semantic Formation in Clusters
To analyze how semantic clusters form in the feature space, we sample 50,000 states per environment and compute: (i) the per-cluster mean image and the mean (± std) pixel distance from each state to its cluster mean (averaged over K=8K{=}8 clusters), and (ii) the probability of cluster transitions along trajectories. We evaluate three models: Trained (ours), Stop-Grad (the ablation in § C.2 that removes the effect of our modules by stopping gradients), and Raw (untrained). Results are shown in Table 4. Compared to Raw, both Trained and Stop-Grad reduce transition probability, indicating policy-induced structure. However, Stop-Grad lacks fully distinct boundaries: it exhibits higher transition probability and lower intra-cluster pixel distance than Trained, whereas our method achieves the lowest transition probability and the highest intra-cluster pixel distance.
Table 4: Cluster transition probability and intra-cluster pixel distance (mean with std) over 50k states.
Environment Model Cluster transition probability Pixel distance mean (Std. Dev.)
@@
Ninja Trained 0.2680 141.57 (67.88)
Ninja Stop-Grad 0.2705 132.46 (66.12)
@@
We further assess temporal coherence with two episode-level metrics: (i) Episode Cluster Entropy (ECE)—the entropy of each episode’s cluster distribution (lower is better, indicating more focused semantic grouping), and (ii) Temporal Cluster Agreement $\text{TCA@}k$—the fraction of frame pairs at lag $k$ assigned to the same cluster (higher is better, indicating smoother, more stable semantics). As summarized in Table 5, our method consistently achieves lower ECE and higher TCA@3/6 than Stop-Grad across all three games. Note that Jumper episodes are shorter, yielding fewer clusters per episode and thus lower ECE values overall.
Table 5: Episode- and frame-wise metrics (averaged over episodes).
Environment Model ECE TCA@3 TCA@6
FruitBot Stop-Grad 1.9068 0.7050 0.5309
FruitBot Ours 1.8886 0.7302 0.5502
Jumper Stop-Grad 0.6987 0.7040 0.6715
Jumper Ours 0.5572 0.7911 0.7089
Ninja Stop-Grad 1.5505 0.5798 0.3923
Ninja Ours 1.3086 0.6787 0.4985
@@ I.6 Part 6: Grouping Details
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and the introduction precisely enumerate three key contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: A dedicated paragraph (§5) details the limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Appendix B states all assumptions and supplies full proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: §4 and App. A list environment names, seeds, architecture, hyper‑parameters (learning rates, loss weights), compute budget, and provide the code in the supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The supplementary zip file contains anonymized source code, configuration files, and scripts that fetch the public Procgen dataset.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Training details (optimiser, batch‑size, PPO clip, etc.) appear in App. A; evaluation follows the Procgen protocol exactly.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All learning curves show mean and error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: App. A states that each run uses a single Tesla V100 (32 GB) for 6 h; total GPU‑days are reported.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work uses only synthetic game data, releases code under MIT license, and poses no identifiable privacy or safety risk.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Appendix J discusses the impacts of this study.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The released artifacts are small game agents that carry no foreseeable dual-use risk.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the Procgen benchmark [19] and note its MIT license in App. A.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No new datasets or models are released beyond trained weights already covered by the code‑release answer.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our user study surveyed adult volunteers on AI models online without studying human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our user study surveyed adult volunteers online, posed minimal risk, and was exempt under our institution’s IRB policy.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: No large language models are used in the proposed methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.